# A Stabilized Unfitted Space-time Finite Element Method for Parabolic Problems on Moving Domains

### v1

@@ 6.4 Boundary layer problem
Interpreting the parabolic equation (2.1) as a convection-diffusion problem in the space-time setting, it is natural to expect similar challenges in convection-dominated regimes. In this subsection, we demonstrate the capability of our method in handling a boundary layer problem, showing that the incorporated SUPG stabilization successfully eliminates numerical oscillations while maintaining accuracy.
We consider a 1D boundary layer problem with the same domain configuration as in Example 6.2. Let $a(x,t)=\varepsilon=2\times 10^{-3}$, $f=1$, $g=u_{0}=0$, and take the background domain as $\widetilde{Q}=(0,1)\times(0,1)$. The initial mesh $\mathcal{T}_{h}$ is generated following the procedure in Example 6.2 with $N_{x}=N_{t}=7$, and is uniformly refined four times to obtain the computational mesh $\mathcal{T}_{h/16}$. Since $\varepsilon$ is smaller than the characteristic mesh size, the problem is convection-dominated.
@@ 7 Conclusions
To ensure the robustness of the proposed method, two stabilization techniques were introduced. First, the ghost penalty stabilization was employed to effectively overcome the small cut problem, thereby controlling the condition number of the stiffness matrix. Second, for handling convection-dominated cases, the streamline upwind Petrov–Galerkin (SUPG) scheme was incorporated to stabilize the temporal advection term.

### v2

@@ 6.4 Boundary layer problem
Interpreting the parabolic equation (2.1) as a convection-diffusion problem in the space-time setting, it is natural to expect similar challenges when $a(x,t)\ll 1$. In this subsection, we demonstrate the capability of our method in handling a boundary layer problem, showing that the incorporated SUPG stabilization successfully eliminates numerical oscillations while maintaining accuracy.
We consider a 1D boundary layer problem with the same domain configuration as in Example 6.2. Let $a(x,t)=\varepsilon=2\times 10^{-3}$, $f=1$, $g=u_{0}=0$, and take the background domain as $\widetilde{Q}=(0,1)\times(0,1)$. The initial mesh $\mathcal{T}_{h}$ is generated following the procedure in Example 6.2 with $N_{x}=N_{t}=7$, and is uniformly refined four times to obtain the computational mesh $\mathcal{T}_{h/16}$. Since $\varepsilon$ is smaller than the characteristic mesh size, the solution exhibits characteristics analogous to those observed in convection-dominated convection-diffusion problems.
@@ 7 Conclusions
To ensure the robustness of the proposed method, two stabilization techniques were introduced. First, the ghost penalty stabilization was employed to effectively overcome the small cut problem, thereby controlling the condition number of the stiffness matrix. Second, the streamline upwind Petrov–Galerkin (SUPG) scheme was incorporated to stabilize the temporal advection term.